### v2

# Measuring the Impact of Missingness in Traffic Stop Data

Abstract
In this article we explore the data available through the Stanford Open Policing Project. The data consist of information on millions of traffic stops across close to 100 different cities and highway patrols. Using a variety of metrics, we identify that the data is not missing completely at random. Furthermore, we develop ways of quantifying and visualizing missingness trends for different variables across the datasets. Because of the way we have identified the missingness dependence on key variables, imputation is not possible and a sensitivity analysis is required. We performing a sensitivity analysis to extend work done on the outcome test as well as to extend work done on sharp bounds on the average treatment effect. We demonstrate that bias calculations can fundamentally shift depending on the assumptions made about the observations for which the race variable has not been recorded. We suggest ways that our missingness sensitivity analysis can be extended to myriad different contexts.
1 Introduction
Every year, more than 20 million traffic stops are conducted in the United States, representing the most common way in which drivers interact with the police (davis2018contacts). Unfortunately, not all traffic stops are conducted equally – in a 2019 survey by the Pew Research Center, 59% of Black men and 31% of Black women say that they have been unfairly stopped by the police (anderson_2020). The relationship between belonging to a minority race, such as Black, Latine, and/or Asian Pacific Islander Desi American, and experiencing discriminatory policing is both historical and happening in the present day. Since the late 90’s, concern over racial profiling has led to federal and state mandates requiring the collection of traffic stop data (russell2001racial).
The Stanford Open Policing Project (SOPP) [^1] https://openpolicing.stanford.edu/, whose data are the focus of our study, has several dozen datasets representing over 100 million separate traffic stops (pierson2020large). [^2] After our manuscript was submitted, SOPP substantially expanded their database in July 2025 and included a Python library for accessing the data https://openpolicedata.readthedocs.io/. In tandem with the increasing availability of policing data is the growing interdisciplinary scholarship analyzing such data for evidence of racial profiling (baumgartner2017racial; grogger2006testing; pierson2020large; smith2001racial; knox2020administrative). The traffic stops literature has developed a range of methods designed to measure racial profiling even in the setting of outcome-dependent sampling, as in the outcome test and the veil of darkness; more recent work frames discriminatory policing as a causal inference problem.
This project re-analyzes two methods of quantifying racially discriminatory policing motivated by concerns of non-ignorable missing data in the original datasets. The motivation is simple: if traffic stop data is meant to hold police officers accountable, but police officers are recording the data, can we really trust the data? To the best of our knowledge, when there is missing data in a covariate or outcome of the model, scholars generally opt to remove those observations as a simple pre-processing step.
Chanin and Welsh’s review of San Diego Police Department traffic stop data (chanin2021) point to the lack of meaningful consideration of data quality in existing traffic stop literature: of the 100 papers they reviewed, the authors find only 19 papers that address missingness and/or recording rates (chanin2021, p. 6). As one may expect, “To an officer who does not see driver race as affecting their own decision-making… data collection only redounds to their detriment” (chanin2021, p. 9).
Motivated by the qualitative evidence of non-ignorable missingness, we conduct a missing data analysis that is organized as follows. Section 2 further contextualizes the missing data problem in the traffic stop setting and defines the different kinds of missingness. Section 3 introduces two metrics for the conditional missingness and numerous visualizations that further diagnose the problem of non-ignorable missingness. Then, we perform two sensitivity analyses in Section 4. The first examines the impact of missing data on the outcome test, a simple descriptive benchmark of the race effect on search from (pierson2020large). The second sensitivity analysis is applied to the causal mediation framework developed by (knox2020administrative) that bounds the Average Treatment Effect (ATE) of race on the use of force during a pedestrian police stop. Our method, summarized in Figure 1, can be applied to datasets and statistics in different domains beyond policing.
Figure 1: A flow chart for our complete sensitivity analysis on the Ohio, Statewide data from SOPP. First, choose the disparity statistic; second, identify which variable(s) have non-ignorable missingness; third, augment the dataset using pseudo-observations; fourth, calculate disparity estimate over all augmentations. See Section 4.1.1 and Section 4.2.1 for details on the process for the outcome test and ATE bounds, respectively.
2 Missingness
The problem of missing data is present across multiple disciplines, like healthcare (kaambwa2012; ward2020; hunt2021; zhang2023machine), natural disasters (jones2022human), psychology (gomer2023realistic), agriculture (robbins2013imputation), epidemiology (allotey2019multiple; balzer2020; tsiampalis2020missing), and crime (edwards2017effect; blasco2021missing; stockton2024now). Some of them derived methods of imputation (robbins2013imputation; blasco2021missing; stockton2024now; bounthavong2015) while others provided innovative approaches to combining information across different missingness mechanisms (ward2020; gomer2023realistic). A few of the papers highlight sensitivity analyses (allotey2019multiple; stockton2024now) to obtain a broader view of the missingness landscape, but their work does not use the missing data in the same way as ours does; in particular, they assume the form of the missingness mechanism (e.g., MAR) and we make no such a priori assumptions.
One particular study by lo2019 on clinical trial data is closely related to our study. The missingness mechanism in the pharmaceutical setting is “largely related to the post-study reporting of clinical trial data” indicating that, as in traffic stop data, there are incentives to obscure the information. By imputing missing data instead of discarding data with missing values, lo2019 achieved superior performance in predicting successful drug approval.
In a pointed summary, meng2024 affirms that the
often fatal problem is that by only learning from the cases where data are complete, one is only learning the patterns of these cases, which are systematically different from those with incomplete data when the reason for incompleteness is related to the outcome we care about.
We expand on the ideas of meng2024 within the context of the SOPP. Part of the motivation for our deep dive into missing data associated with traffic stops is due to policies and laws that require officers to guess at the age, race, and other demographic characteristics of the driver who has been stopped. Guessing at someone’s age or race (and sometimes choosing not to guess) creates complete cases which are systematically different from the incomplete cases (e.g., see Figure 2).
Figure 2 shows two examples of some of the unusual structures in the data: in North Carolina, the age of stopped drivers has peaks in increments of five, which probably reflects some policy of officers recording perceived age instead of actual age. In Seattle, the race for stopped drivers is almost always unrecorded. Note that the lack of race data makes the Seattle SOPP dataset inappropriate for an analysis on racial profiling in traffic stops. The issues in North Carolina and Seattle motivate us to think carefully about using the SOPP for large scale analyses on racial profiling in traffic stops. The problems in the dataset may result from data corruption along the way, or alternatively, they may now be resolved in more current downloads of the data. However, like many others, we use the SOPP, and the two examples foreshadow the challenge of accurately comparing racial bias across datasets when the data collection schemes are so different. In Section 4, our sensitivity analyses are limited to only datasets where missing race values correspond to non-missing information in the search and contraband found variables.
Figure 2: Examples of idiosyncratic data collection from the SOPP database in Charlotte, North Carolina (left) and Seattle, Washington (right). In Charlotte, the age distribution of stopped motorists has peaks at increments of five. In Seattle, the race of motorists is almost never recorded.
Consider the following quotes from three different assembly bills (boldface ours).
  • CA Assembly Bill No. 953 (CA953),
    “The perceived race or ethnicity, gender, and approximate age of the person stopped, provided that the identification of these characteristics shall be based on the observation and perception of the peace officer making the stop, and the information shall not be requested from the person stopped. For motor vehicle stops, this paragraph only applies to the driver, unless any actions specified under paragraph (7) apply in relation to a passenger, in which case the characteristics specified in this paragraph shall also be reported for him or her.”
  • New York Assembly Bill 3949 (NY3949)
    “…each law enforcement agency shall, using a form to be determined by the division, record and retain the following information with respect to law enforcement officers employed by such agency…the characteristics of race, color, ethnicity, gender and age of each such person, provided the identification of such characteristics shall be based on the observation and perception of the officer responsible for reporting the stop and the information shall not be required to be provided by the person stopped;”
  • Tennessee House Bill 2167 (TN2167)
    “Characteristics of race, color, ethnicity, gender, religion, and age of any person stopped for a traffic violation. The identification of such characteristics must be based on the observation and perception of the law enforcement officer responsible for reporting the stop, and the person stopped will not be required to provide the information.”
One notable trend of these policies is that usually the race, gender identity, and age is recorded based on officers’ perception. The Stanford Open Policing Project’s own analysis of its data points out this detail (pierson2020large). We can see evidence of “perceived age” in how the age density plots of North Carolina datasets spike at multiples of five (see Figure 2), but we can also see evidence of officer perceptions in the extensive exclusion of both non-binary gender identities and mixed-race identities and the common inclusion of Hispanic, which is an ethnicity, as a level for race. Another important feature of such state mandates are the different variables required for each state. To assess larger trends of missingness, we are able to compare data quality and missing data across datasets with myriad different variables, but datasets that collect neither search nor contraband found (or other acts of force) are outside the scope of most statistical models testing for discriminatory policing.
Our work differentiates two distinct types of missingness: (1) missing data and (2) missing potential outcomes. Missing data occur when observed data are incomplete, with values unrecorded. Missing data can be categorized as Missing Completely at Random (MCAR), Missing at Random (MAR), or Missing Not at Random (MNAR); complete definitions of the terms are given in Section 2.1. In contrast, missing potential outcomes arise inherently within the framework of causal inference, where for each unit, only the outcome corresponding to the treatment received is observed, while counterfactual outcomes remain unobserved.
In a comprehensive assessment of the current missingness literature, mitra2023 point out how “literature on causal inference with missing data is very scarce, and it is particularly challenging when both input variables and potential outcomes are systematically missing.” The primary focus of this project is about non-MCAR missingness – diagnosing it and observing how it can impact downstream analyses including causal inference methods. In order to separate these two forms of missingness clearly, we define the following terms.
2.1 Definitions of missingness
2.1.1 Missing data
Missing data refers to an incomplete observed dataset, i.e., data that was intended to have been collected but was not. Missing data is common in all types of data. In the context of policing, because much of the data are self-reported, it is not surprising that there would be extensive missing data. We refer the reader to meng2000, who gives an excellent tour through details of missingness, nonresponse bias, and modeling of real-life missing data mechanisms. In order to describe the missing data mechanisms, we rely on the characterizations from little2020:
  • MCAR
    If missing data do not depend on the values of the data, missing or observed, that is, the probability of missing is the same for all levels of each variable, the data are called missing completely at random (MCAR) (little2020). Note that the definition implies that the missing data cannot be related to the research question, here, the race effect.
  • MAR
    If the missing data depend on a particular variable only through the observed components of that variable, the missingness mechanism is then called missing at random (MAR) (little2020). Usually, the missingness associated with MAR data has a probability of being missing which is a function of other variables in the dataset which are completely observed.
  • MNAR
    We refer to data as Missing Not at Random (MNAR) if the distribution of the missingness depends on the value of the missing components themselves (little2020) or an external unrecorded variable. The probability of data being missing is related to what the missing values would have been, had we observed them. More recent work has extended MNAR to include any missing data that is not MCAR or MAR, that is, when the function describing the probability of missing data is unknown (lo2019).
We note that MCAR and MAR are ignorable missingness. Estimates computed on MCAR/MAR data will typically be more variable (due to smaller sample sizes), but no more biased than estimates computed on complete data. Additionally, ignorable missingness allows for the possibility of performing analyses using only the observed data without introducing bias, but they require appropriate methods for analysis. The missingness due to MNAR is non-ignorable, as estimates will be unreliable and potentially quite biased. We emphasize that “without introducing bias” is the key to ignorable missingness (without the phrase, the idea is meaningless because one could always perform an analysis). Importantly, in order for the missing data mechanism to be truly ignorable, the parameter(s) in the model describing the data should have no tie(s) to the parameter(s) that regulate(s) the missing data mechanism (rubin1976).
2.1.2 Missing potential outcomes
While much work has been done to estimate the race effect in traffic stops (see Section 1), the task of identifying race as a causal mechanisms is much harder than estimating the race effect. While certainly race is an individual characteristic, we follow knox2020administrative who consider race to be an aspect of the police-driver encounter rather than an aspect of the driver’s identity. They say, “The manipulation of race is conceptualized as the counterfactual substitution of an individual with a different racial identity into the encounter, while holding the encounter’s objective context—location, time of day, criminal activity, etc.—fixed. In other words, the ‘treatment’ in this case is the entire ‘bundle of sticks’ encapsulating the race of the civilian—including, for example, skin tone, dialect, and clothing” (knox2020administrative).
In Figure 3, we extend the Directed Acyclic Graph (DAG) from knox2020administrative to illustrate the causal ordering of variables (and missing potential outcomes) in the SOPP datasets along with the missing data aspects of the variable relationships (indicated by $R$ in the graph); the knox2020administrative DAG is provided by the red lines in Figure 3 (note that they include race in $X_{1}$ because their assumption is that all variables are complete). In their work, knox2020administrative assume that any missing data (across all variables) is MCAR. One of their key assumptions is “Mandatory Reporting” explained as “encounters that escalate to the use of force also trigger a reporting requirement and are, therefore, observed in administrative data” (knox2020administrative). Indeed, the work of knox2020administrative is about the counterfactual of not having been stopped, a missing potential outcome (and not about missing data). We challenge their assumption (of MCAR) and build on their work (of missing potential outcomes) to incorporate data which is missing not completely at random.
Based on the ideas of moreno2018canonical (missing data) and knox2020administrative (missing potential outcomes), we create a DAG model which incorporates both missing data and missing potential outcomes in the context of traffic stops (see Figue 3). The variables and their causal ordering are justified in more detail in Appendix B (in particular, Table 6 describes each connecting arrow), but importantly, in our DAG, race is a variable in the set $X_{2}$ (covariates with missing data) which can causally determine not only the traffic stop (mediator $M$) and the outcome ($Y$) (e.g., being searched) but also the missing data aspect of all three variables (e.g., race ($R_{X_{2}}$), stop ($R_{M}$), search ($R_{Y}$)).
The arrows from $X_{2}$ (missing data covariates) to the missingness indicators $R_{X_{2}}$ suggest that the missingness mechanism is MNAR (as described in Section 2.1). Additionally, the DAG in Figure 3 allows for missing data to be affected by the value of the outcome $Y$ (e.g., whether an individual was searched affects missingness in records of their stop). The DAG in Figure 3 is similar to DAG (J) in moreno2018canonical who conclude that it is not possible to recover the relationship between the outcome and the treatment in such an MNAR scenario. They say (our work is analogous to their m-DAG J):
Both m-DAG E and m-DAG J appear plausible. In both cases, the proportion exposed is nonrecoverable and sensitivity analyses would be required… With m-DAG E, the regression-adjusted exposure-outcome association can be unbiasedly estimated using common methods, but sensitivity analyses would be required with m-DAG J.
Because the relationship is unrecoverable, we use a sensitivity analysis to provide an estimate of the range of possible values of our estimates of interest (see Section 4.2).
Figure 3: Extending the DAG in knox2020administrative to incorporate missing data. Bold arrows depict causal mechanisms that lead to missingness. The presented graph is similar to DAG (J) in moreno2018canonical suggesting nonrecoverability and necessitating sensitivity analysis. $X_{1}$ represents completely observed covariates like time and day. $X_{2}$ represents covariates with missing information, like age and race. The knox2020administrative DAG is provided by the red lines.
3 Diagnosing non-MCAR
3.1 About the data
The SOPP consists of 88 datasets (see Table LABEL:tab:opp-data) of pedestrian and vehicle stops. Each police department and highway patrol has its own idiosyncratic data collection - the years of data collection, number of observations, and number of variables are all different. To facilitate comparisons across datasets that record different numbers of variables, we narrow down our analysis to only include twenty of the most commonly recorded variables among the 88 datasets. The variables are listed in Table 4 in the Appendix. Using a subset of variables ameliorates unfair comparisons across datasets – for example, a dataset with large amounts of missing values for vehicle color is not as problematic as a dataset with large amounts of missing values in subject race. For each dataset, we only analyze variables from the list of twenty. Domain knowledge provides another way to narrow down the list of relevant variables.
3.2 Measuring missingness
We begin with a toy example of a conditional missingness distribution. Consider the following five rows from vehicular stops in Chicago, Illinois on March 14th, 2016 in Table 1. Assume only five traffic stops occurred that day; also assume there are four variables recorded: date, time, race, and age.
date time subject_race subject_age
2016-03-14 08:00:00 white NA
2016-03-14 20:00:00 black 18
2016-03-14 15:00:00 white 24
2016-03-14 17:00:00 NA NA
2016-03-14 20:00:00 hispanic 48
Table 1: Toy example of five observations from the Chicago, Illinois dataset.
By glancing at each column of the table, we can quickly compute the amount of missingness in each variable when conditioning on the date being March 14, 2016: the variable time has no missingness; subject race has a 20% missingness rate; and subject age has a 40% missingness race on that particular day. Then, we could repeat these calculations for each unique value of date in the dataset, and we would recover the empirical date-conditional missingness distributions, one for each of the three variables time, race, and age.
Moreover, we could aggregate the three conditional missingness distribution to produce a dataset-level conditional missingness distribution on date. In Table 1, the dataset-level missingness on March 14, 2016 is 20% (the average of each variable-specific missingness percentage.)
The introductory example can be formalized as follows. We first identified a conditioning variable (date). Next, we partitioned the conditioning variable into equal-sized bins by day, although we could have chosen something like week or month. Finally, we computed two types of empirical conditional missingness distributions. The first type was specific to the variable: we looked at the percent of NA values in each variable and bin. The second type provides a dataset-level measure of the missingness by averaging the variable-specific empirical conditional missingness distribution. Now, we formalize the strategy into notation.
Let $X$ denote one of the 88 traffic stop datasets with $n$ observations and $m\leq 20$ variables. [^3] There are 88 datasets, so perhaps one could write $X^{(d)}$ for $1\leq d\leq 88$. But the definitions only apply to one dataset at a time, so we avoid the extra notation. We will denote $X_{ij}$ as the $(i,j)$ entry of the dataset; $X_{\cdot j}$ as the $j$-th column; and $X_{i\cdot}$ as the $i$-th row. The binary missingness matrix $\omega$ is defined entry-wise as $\omega_{ij}=\mathds{1}\left\{X_{ij}=\texttt{NA}\right\}$.
The distributions of $\omega_{ij}$ are unknown, so it is impossible to ask if $\omega\perp X$, or MCAR, holds. Instead, we study the conditional missingness $\omega_{ij}\mid X_{ik}$, where $X_{\cdot j}$ may be some sensitive attribute and $X_{\cdot k}$ is the conditioning variable, a neutral and continuous variable with naturally low levels of missingness (e.g., date). Our strategy cannot be directly used to test for MNAR. Under MNAR, the conditional missingness is dependent on something unobserved, such as the missing data itself. For the traffic stop data and other real world datasets, natural choices of the conditioning variable $k$ include time, date, latitude, and longitude, all of which usually have low levels of missingness and are automatically recorded.
After selecting a conditioning variable, we partition the conditioning variable into $B$ ordered bins, which we denote as $b\in\{1,2,\ldots,B\}$. We write the corresponding index set of a bin $b$ as $I_{b}$. There is usually a natural choice of bin – if using date as a conditioning variable, $b$ can be on the units of weeks or days; if using time as a conditioning variable, $b$ can be on the unit of hour. Next, for each $b$, we can compute the percent of missingness for observations in that bin $I_{b}$. The percent of missingness can be calculated either for a single sensitive attribute $X_{\cdot j}$ (e.g., the percent of missingness in race on a particular day) or across all m1m-1 variables excluding the conditioning variable (e.g., the percent of missingness on a particular day, aggregated across all variables).
We define these two measurements of conditional missingness. The first is for a single covariate; the second aggregates over m1m-1 covariates. [^4] The m1m-1 in the denominator is because the missingness of the conditioning variable itself is not considered.
  • The Conditional Missingness Rate (CMR) of covariate $j$ given some conditioning variable $k$ taking values in bin $b$ is the average amount of missingness in covariate $k$ for observations in bin $b$.
    $\overline{\omega}_{j}(k,b)=\frac{1}{|I_{b}|}\sum_{i\in I_{b}}\omega_{ij}\qquad\text{where}\qquad I_{b}=\{i\in[n]\mid X_{ik}\in b\}.$ (1)
  • The dataset Conditional Missingness Rate (dCMR) given some conditioning variable $k$ taking values in bin $b$ is the average amount of missingness in the entire dataset for observations in bin $b$.
    $\overline{\omega}(k,b)=\frac{1}{m-1}\sum_{j\neq k}\overline{\omega}_{j}(k,b).$ (2)
The purpose of dCMR is to see the relationship between missing values and some conditioning variable. While we do not formally test the independence of the conditioning variable and the missingness (as one might do if they were formally testing for non-MCAR), we will later compute the maximal correlation between the conditioning variable and the dCMRs. In the plots below, the calculation of dCMR is done by averaging over the set of variables (when available) given in Table 4 for each dataset.
The maximal correlation measures non-parametric dependence between two variables (breiman1985estimating). Maximal correlation does not assume a linear or monotonic relationship between two variables (unlike the Pearson correlation or Spearman’s rank correlation, respectively). The maximal correlation finds transformations of each variable that maximizes the population-level correlation. While it can be used to formally test for independence, we use maximal correlation as a summary statistic to measure the dependence of two variables. A high value of maximal correlation indicates that there is a (nonlinear) relationship between the dCMR and the variable of interest (see, for example, Figures 4 - 8 where all of the relationships have quite high maximal correlation). We are not able to use standard tests of MAR (e.g., Little’s Test (little1988)) because our data are not normally distributed. The maximal correlation gives us a sense of dependency between missingness and other variables without using a formal test. We use the R package acepack which implements the Alternating Conditional Expectations algorithm to compute the maximal correlation (spector2016package).
When auditing a large number of datasets and variables for non-MCAR, the maximal correlation offers a quick way to understand the missingness patterns at a high level. If most datasets have low maximal correlation, then an analyst might not need to manually check each dataset for missingness patterns.
3.3 Notable dCMR trends across the SOPP
Here we provide a high-level audit of the entire 88 datasets (see Table LABEL:tab:opp-data) in the SOPP when conditioning on date and binned by week. Date as a conditioning variable is especially natural because all 88 datasets record date (although some datasets have some missingness in date).
We reviewed the dCMRs for all datasets and selected the six plots in Figure 4 as representative of common trends within the SOPP datasets (all of which seem to be be not MCAR). The problematic relationships between dCMR and date (by week) include piecewise constant; changes in variance; and non-monotonic trends across time. The information within the data itself does not provide reasons for the patterns; they could be due to procedural changes, changes in how the data were collected, or changes in missingness for particular subsets of variables. If the research goal is to truly understand the exact missingness patterns, we recommend a further dive into the CMR plots for individual variables. Note, however, that our work has found no strong trends between the additional variables and the variables of interest in our work (i.e., race and outcome measures like searched and contraband found). The dCMR (missingness on the y-axis) for each of the plots is on a different scale. We let the scales be free in order to focus on the trends in the missingness patterns.
Figure 4: The dCMR of six selected datasets conditioned on the date (binned into weeks). The number given in parentheses is the maximal correlation between dCMR and date. The datasets demonstrate heterogeneity over time: level changes, changes in variance, and non-monotonic changes.
In Figure 5 we demonstrate the usefulness of the maximal correlation by selecting the five plots corresponding to the minimum maximal correlation, the maximal correlation quartiles (25th, 50th, and 75th percentiles), and the maximum maximal correlation, providing a high level summary of the extent of dependence between missingness and date. One might argue that we cherry-picked the datasets in Figure 4. However, the maximal correlation values in Figure 4 are around the $50^{th}$ and $75^{th}$ quantiles of the maximal correlation distribution (seen in Figure 5), indicating that the datasets described in Figure 4 are not extreme cases compared to the overall trends we see across all of the datasets. Note that the dCMR (missingness on the y-axis) for each of the plots is on a different scale. We let the scales be free in order to focus on the trends in the missingness patterns.
Figure 5: The dCMR of five datasets conditional on date (binned into weeks) corresponding to the minimum, quartiles (25th, 50th, and 75th percentiles), and maximum maximal correlation arranged from left to right; in parentheses is the maximal correlation between dCMR and date.
3.4 Auditing eight datasets for non-MCAR
In Section 4, we consider how non-MCAR bias can be assessed in models of discrimination that rely on three covariates: race, whether or not a search was conducted, and whether or not contraband was found. Thus, here, we tailor our analysis to the eight datasets that have at least 10% missingness in at least one of race, search, or contraband (with all three of the variables recorded for the dataset). The eight datasets consist of seven statewide datasets and only one city-level dataset: California; Colorado; Chicago, Illinois; Maryland; New Jersey; Ohio; Washington; and Wisconsin. Of the eight datasets, eight record the variable date; six record the variable time; and four record the variables latitude and longitude. The dCMR is computed using the 20 variables described in Table 4.
3.4.1 Temporal trends
First, we examine how missingness correlates with date in Figure 6. Although the scale of the y-axis is small for many of these plots, we clearly see non-random missingness trends over time. Three datasets (California, Colorado, and Washington) have increasing missingness over time; two (Chicago and New Jersey) have decreasing; and three (Maryland, Ohio, and Wisconsin) have trends that are roughly equal over time, with each state having a handful of years that are clearly outliers with respect to the missingness. As with Figure 4, we cannot know the reason for the missingness patterns. A deeper dive into the CMR plots for individual variables could shed light into some of the relationships, however, we have found that none of the variables are strongly associated with the missingness trends connecting race and outcome variables (like search and contraband found).
Figure 6: The dCMR of eight datasets conditional on date (binned into weeks). The maximal correlation for each dataset is included in parentheses.
When considered across time (binned by hour), the dCMRs also exhibit dependence with the conditioning variable in Figure 7. In four datasets (Chicago, New Jersey, Ohio, and Wisconsin), missingness peaks in the early mornings; in two datasets (Maryland and Washington state), missingness has two local peaks – one in the early morning, and one later in the day. However, we point out that the range of dCMR missingness is quite small, and the peaks in missingness may have to do with external factors.
Figure 7: The dCMR of six datasets given the time (binned into hours). The maximal correlation for each dataset is included in parantheses.
3.4.2 Spatial trends
When we condition on latitude and longitude, we can plot each bin (6-digit geohash) with the $x$-axis as longitude and the $y$-axis as latitude. Then, the dCMR can be color-coded. For the three state-level datasets in Figure 8, the dCMR plot on latitude and longitude naturally recovers state highways (although it is not a perfect representation due to some highway patrols working primarily on highways (WA) and some patrolling side streets as well (OH)). The maximal correlation here is an average of the maximal correlation between latitude and dCMR and between longitude and dCMR. The maximal correlation between location and missingness in state-wide datasets is low, with only Wisconsin state having maximal correlation above 0.1. Near city centers like Minneapolis towards the west and Milwaukee in the east, the dCMR is slightly higher than its surroundings. Meanwhile, in Chicago, one can clearly see the higher levels of missingness in the city center and lower missingness in the suburbs.
Figure 8: The dCMR of four datasets given the latitude and longitude (binned into 6-digit geohashes). The maximal correlation (average of the maximal correlation of latitude & dCMR and longitude & dCMR) for each dataset is included in parentheses.
4 Assessing sensitivity of missingness
Using the SOPP, we analyze several datasets to broadly explore differential missingness. As a way to investigate the impact of the missing race data, we assign observations whose race is not measured to either the Black or white groups of individuals, randomly. Subsequently, we apply two published methods to the augmented dataset in order to investigate how missingness can impact racial bias estimates. Although both published studies consider race within the framework of missing potential outcomes, neither study delves into the impact of missing data in considering the reporting of a race effect. Our sensitivity analysis furthers the literature on understanding racial discrimination in traffic policing. Our work has similarities to meyer2024 who use a sensitivity analysis on a new test, the Overlapping Condition Test.
4.1 The outcome test
An analysis of over 20 million traffic stops applies the outcome test (originally described in becker1993nobel applied to traffic data in pierson2020large) to compare the rate of searches to the rate of successful searches in order to determine if the police officers use a lower standard for searching drivers of color (race effect). The researchers find evidence of both racial profiling and preferential searches for minority race drivers. Although the discussion in pierson2020large focuses on standardizing and improving the quality of traffic stop data collection, it does not mention how missing data plays a role in the analyses.
Applied to the traffic data, the outcome test (pierson2020large, Figure 3) compares “hit rates” (point estimates of the probability of recovering contraband given that a driver is searched) between Black and white drivers. The idea is that if searches of minority drivers reveal contraband at lower rates than searches of white drivers, the outcome test suggests that officers are searching minority drivers on the basis of less evidence and with lower success, i.e., they are potentially racially biased.
The outcome test is not designed to identify causal mechanisms for differential hit rates; instead, it is a measure of the race effect. For example, a lower hit rate could be due to racial bias of the officers or because Black drivers are carrying around less contraband than white drivers. However, the latter reason would require that (1) the police be colorblind and (2) the police are not using any other correlated indicators related to contraband (e.g., the age of the vehicle). Our work does not delve into the shortcomings of the outcome test (e.g., inability to identify causal indicators), which have been studied in detail in prior work (engel2008critique). Instead, we offer a sensitivity analysis that assesses the impact of missing data race information. We offer our sensitivity analysis to complement analyses such as becker1993nobel, persico2006generalising, pierson2020large, and gaebler2025 who use the outcome test in their own work.
4.1.1 Sensitivity analysis of the outcome test
Throughout our sensitivity analysis, we use disparity as a measure of racial discrimination. Equation (3) provides the generic form of disparity where $p$ represents the probability of some police action (e.g., being searched). Equations (4) and (5) are specific to the outcome test (here, Section 4.1.1) and bounds on the ATE (see Section 4.2.1), respectively. It is important to note, however, that $\text{disparity}_{\text{outcome}}<0$ indicates discrimination against Black drivers because if Black drivers have lower rates of contraband found than white drivers, they are being searched unfairly. On the other hand, $\text{disparity}_{\text{ATE}}>0$ indicates discrimination against Black drivers because the traffic stop “as if Black” is being searched more often than the traffic stop “as if white.”
disparity $=$ $p_{Black}-p_{white}$ (3)
$\text{disparity}_{\text{outcome}}$ $=$ $P(\text{contraband }|\text{ stop, search},\text{Black})-P(\text{contraband }|\text{ stop, search},\text{white})$ (4)
[…]
The first step of the sensitivity analysis is to calculate the ignore NA disparity: point estimate of the probability of recovering contraband given that a driver is searched for Black drivers minus white drivers (without considering any observations with NA race).
Next, we ask how the disparity would change if the observations with missing race had not been missing. For example, what would the disparity be if all of the observations with missing race turned out to be Black drivers? What would the disparity be if all the observations with missing race turned out to be white drivers? Or something in between? As described below (and worked out in detail in Table 3), our sensitivity analyses create new datasets (referred to as augmented data) by allocating the NA race observations to either the Black or white group. We call the allocated NA race observations pseudo-observations; we use the word “pseudo” because they represent actual observations from the dataset, but their race has been augmented through the sensitivity analysis.
Figure 9 presents a sensitivity analysis based on allocating NA race drivers to either the Black or white driver groups for a random subset of counties in Ohio (the image for all Ohio counties is given in Appendix C in Figure 15). The left plot of Figure 9 presents the ignore NA county-level disparities. Note that all of the ignore NA county-level disparities are greater than zero, which indicates that there is no evidence of discrimination against Black drivers. The right plot presents all possible disparities from assigning NA race observations as either Black or white. That is, every possible allocation (see Table 3) of the missing race observations across the Black and white groups is considered, and the disparity is calculated on the augmented data. Each dot in the right plot represents one possible allocation of the NA race observations. We recognize that allocating all of the NA race drivers to Black can happen in exactly one way; however, allocating half of the NA race drivers to Black can happen in many ways (assuming each individual driver is distinct). We calculated the number of combinations, not the number of permutations because using the frequency associated with the number of permutations would imply that the NA race drivers were equally likely to be distributed across Black and white drivers. Instead, the sensitivity analysis allows us to see the possible values given any potential distribution of the NA drivers to Black and white. In the right plot, the ignore NA disparity estimators are given as open black circles, which trend more positive (less discrimination against Black drivers) than the vast majority of measurements obtained from the sensitivity analysis (possible disparities across all potential allocations of NA drivers to Black and white). Tables 2 and 3 detail the process of all possible allocations of NA drivers to the Black and white driver groups for Belmont County, OH (visualized in the middle of Figure 9).
We see that incorporating missingness leads to estimates with lower disparities, indicating that incorporating the missing information tells us that more discrimination against Black drivers is happening than we had previously understood (with the ignore NA estimator only). The color of the point in the right panel of Figure 9 indicates the proportion of missing race observations that were allocated to the white group for calculations.
Table 2: Original dataset for Belmont County, OH.
Contraband found No contraband found Total Naive hit rate
Black 45 170 215 45/215 = 0.209
white 286 1726 2012 286/2012 = 0.142
NA 4 643 647 4/647 = 0.006
Table 3: Augment the dataset by allocating NA race observations to Black and white groups for Belmont County, OH.
NA race
Contraband
$\to$ B
NA race
Contraband
$\to$ W
NA race
No contraband
$\to$ B
NA race
No contraband
$\to$ W
disparity
Equation (3)
Sums to 4 Sums to 643
4 0 643 0 $\frac{45+4}{215+4+643}-\frac{286+0}{2012+0+0}=-0.085$
3 1 643 0 $\frac{45+3}{215+3+643}-\frac{286+1}{2012+1+0}=-0.087$
0 4 643 0 $\frac{45+0}{215+0+643}-\frac{286+4}{2012+4+0}=-0.091$
4 0 0 643 $\frac{45+4}{215+4+0}-\frac{286+0}{2012+0+643}=+0.116$
0 4 0 643 $\frac{45+0}{215+0+0}-\frac{286+4}{2012+4+643}=+0.100$
To quantify the changes from the ignore NA estimator, note the following, which are summarized in Table 4.1.1. Of the 87 counties in Ohio, 10 indicate discrimination against Black drivers (10 open dots to the left of the vertical dashed line at zero in Figure 15) and 77 indicate discrimination against white (77 open dots to the right of the vertical dashed line at zero in Figure 15). However, when considering the sensitivity, 71 counties could switch disparity estimate (71 points could switch sides of the vertical line at $x=0$): 66 could switch from discrimination against white to discrimination against Black drivers, 5 could switch from discrimination against Black drivers to discrimination against white drivers, and 16 counties could not switch discrimination, as measured by the sensitivity analysis.
[!htb]Outcome test disparity estimates. $d=p_{Black}-p_{white}>0$ (++) indicating discrimination against white drivers; $d=p_{Black}-p_{white}<0$ (-) indicating discrimination against Black drivers. Dataset Total number of counties/ departments Counties/ departments with missingness Sign of ignore NA disparity estimate Imputed county-level disparity estimate (-) (++) Counties that could switch signs Counties that don’t switch signs (-) to (++) (++) to (-) Remains (-) Remains (++) OH 89 87 10/87 77/87 5 66 5 11 CO 66 32 25/32 7/32 0 2 25 5 WI 72 49 25/49 24/49 3 2 22 22 WA 39 31 19/31 12/31 3 3 16 9 MD 223 60 34/60 26/60 0 2 34 24
Appendix C contains the same sensitivity analysis for the states described in Section 3.4. Note that we previously identified eight datasets that had missingness of at least 10% in each of the variables race, search, and contraband. However, New Jersey and California plus Chicago, IL datasets do not contain relevant information for the outcome test. That is, when race is NA, all of search conducted and contraband found are also NA in NJ, CA, and Chicago. We can get ignore NA estimators for each state, but we cannot repeat our sensitivity analysis for those states where missingness in race, search, and contraband happens simultaneously. The analysis of missing race data is given in full for Ohio, Colorado, Wisconsin, Washington, and Maryland highway patrols in Appendix C.
Figure 9: On both images, the ignore NA county-level disparities are given by open black circles sized by the number of total searches from that county in Ohio. The right plot presents all possible disparities from assigning NA race observations to either Black or white. Boxplots describe the distribution of augmented disparities across the allocations of missing race observations. In each boxplot, the black horizontal line is given as the median of the augmented disparities across all possible allocations. The color of the point in the right panel indicates the proportion of missing race observations that were allocated to the white group.
4.2 Bounds on the average treatment effect (ATE)
knox2020administrative conceptualize race as a causal variable (missing potential outcomes) and the “investigation of racial bias in policing as an inherently causal inquiry.” Using causal inference, they derive strict bounds for the Average Treatment Effect (ATE) (the extent to which drivers of color face greater risk of force than white drivers because of their race), given a traffic stop was made. In particular, the ATE measures the probability of being searched if a person was Black and stopped as if they were Black minus the probability of being searched if a person was white and stopped as if they were white, given that a stop was made, see Equation 5. They do not discuss missing data or its impact on the analyses. However, they do require the “Mandatory Reporting” assumption (that all encounters are recorded) in their derivations (an assumption which is clearly violated if some data are missing).
knox2020administrative describe the missing potential outcomes seen in Figure 3 as the triangle between $X_{2}$, $M$, and $Y$. We extend their work to include missing data ($R_{X_{2}}$, $R_{X}$, and $R_{Y}$). Because it is not possible to recover the relationship between outcome and treatment in the missing data / missing potential outcomes scenario (moreno2018canonical), we provide simulated bounds to assess the impact of $R_{X_{2}}$, missing data.
knox2020administrative derive nonparametric bounds of ATE (given a traffic stop was made) on being searched, as a function of $\rho$, the proportion of racially discriminatory stops. Note that $\rho$ is unknowable, and the knox2020administrative results are provided over a range of possible $\rho$ values. As such, they are able to measure the missing potential outcomes information (in the complete data case), but they do not evaluate the impact of missing data on the missing potential outcomes analysis.
4.2.1 Sensitivity analysis of the bounds on average treatment effect (ATE)
To address possible non-MCAR missingness in race on ATE, we perform another sensitivity analysis where we calculate the ATE bounds when NA values have been ignored (equivalent to knox2020administrative which are based on bounds by horowitz2000nonparametric) and compare with the bounds when NA values have been allocated to the Black and white groups. Consider Colorado and Washington highway patrols, as seen in Figure 10. Following knox2020administrative, by ignoring the NA race values, we calculate the naïve disparity and also bounds on the disparity for each of three different values of $\rho$ (the proportion of racially discriminatory stops): 0.25, 0.5, 0.75. For each of three different values of $\rho$ we calculate the bounds on the ATE, while ignoring the NA values (just as knox2020administrative). Those bounds are represented by the dashed vertical lines and the corresponding ribbon plots. The solid vertical lines are the Knox et al. ATE bounds at each of a different value for the proportion NA assigned to white. The red dots represent the estimated ATE (naïve estimate) at each value for the proportion NA assigned to white. As expected, a higher proportion of racially discriminatory stops naturally leads to a higher range of values of disparity.
The sensitivity of the ATE disparity estimator did not enumerate all possible combinations associated with allocating the NA race observations. Instead, to create pseudo-observations, we allocated a fixed proportion of NA race observations to white (0, 0.25, 0.50, 0.75, and 1.0), and, for each proportion, we randomly selected that proportion of the NA race drivers (and their search outcome) to assign to the white drivers and the remaining NA race drivers (and their search outcome) were assigned to the Black drivers, resulting in an augmented dataset. Additionally, we did two extreme sensitivity analyses: first, we assigned all of the NA race and searched drivers to the Black driver group (and all of the NA race and non-searched drivers to the white driver group); second, we assigned all of the NA race and searched drivers to the white driver group (and all of the NA race and non-searched drivers to the Black driver group).
Recall that $\text{disparity}_{\text{ATE}}>0$ indicates discrimination against Black drivers. Figure 10 (and additional figures in Appendix D) show that in most original estimates (i.e., not including NA drivers) and sensitivity analyses, the disparity indicates discrimination against Black drivers.
In many of the datasets, there are very low rates of search = TRUE within the NA race drivers. Additionally, the group of Black drivers is usually much smaller than the group of white drivers. Therefore, when assigning most of the NA race drivers to the Black driver group (i.e., proportion NA assigned to white is low), the Black driver group seems less frequently searched than ignore NA estimates (i.e., the discrimination against Black drivers seems lower when most of the NA race drivers are assigned to white).
Note that the “extreme” cases (Figure 10 and Appendix D) seem to violate the strict bounds of knox2020administrative. However, the bounds from knox2020administrative are only strict if there is no missingness (or if the missingness is MCAR) and are derived given the DAG of knox2020administrative which is the red subset of the DAG in Figure 3. In Figure 3 and Section 3 we justified arrows from the variables to the missingness indicators which motivate our sensitivity analysis. And the sensitivity analysis demonstrates that the strict bounds of knox2020administrative are not strict if the NA race observations are considered.
Appendix D contains the same sensitivity analysis as Figure 10 (proportion of NA race drivers assigned to white as well as extreme value plots) for the states described in Section 3.4. Note that we previously identified eight states that had missingness of at least 10% in one of the variables race, search, and contraband. As before, New Jersey and California highway patrol and Chicago, IL datasets do not contain relevant information for the ATE sensitivity analysis. That is, when race is NA, all of search conducted is NA in NJ, CA, and Chicago. We can get estimates that ignore NA, but we cannot perform the sensitivity analysis.
Figure 10: Dashed line at zero indicates no differential racial impact. Disparity is the difference in proportion of search for Black minus proportion of search for white. Disparity is measured as bounds on the search ATE, given recorded outcome, for different levels of the $\rho$ parameter (the overall proportion of racially discriminatory stops). The $x$-axis of each plot presents a different allocation of assigning NA race observations to either Black or white.
4.3 Discussion of sensitivity analysis
The goal of our work was to investigate the impact of including NA race drivers in methods used to identify racial discrimination of traffic stops. That is, is the race of the driver being covered up to avoid any perception of discrimination? We were able to show that across the SOPP the datasets have quite a bit of missing information that is not MCAR. By focusing on observations that have missing race data, we were able to perform sensitivity analyses to understand the impact of missing race on two different discrimination metrics.
The outcome test looks at discrimination in rates of contraband found. Table 4.1.1 demonstrates that the NA race drivers, when added to the non-NA race drivers, can change our perspective on discrimination. In particular, Ohio shows that in most counties, it is possible for the NA race drivers to change the perceptions from discrimination against white drivers to discrimination against Black drivers. Other counties show some of the same effect but to a lesser degree.
The sensitivity analysis done on the metric which provides strict bounds on the ATE gave insight into discrimination based on the search variable. The ignore NA statistics indicated that there is some discrimination against Black drivers with respect to the search variable. In Ohio, the discrimination still exists, even in the most extreme case of the sensitivity analysis. While some other states do show that there could be a lessening of discrimination against Black drivers (when NA race drivers are included), the general sense is that, due to the sensitivity analysis, there is anti-Black discrimination as measured by the ATE disparity.
5 Discussion
As a way to objectively assess the amount of missingness in a dataset, we define two metrics in Section 3.2 to quantify the distribution of missingness given some conditioning variable. Our metrics are able to identify both temporal and spatial relationships, and we point out that missingness patterns across either time or space could indicate problematic data collection habits of the peace officer.
Our work demonstrates how missing data information can be used to understand missingness trends and the sensitivity of racial disparity outcomes to missing data. We clarify the distinction between missing data and missing potential outcomes and reason that understanding both is important. Our sensitivity analysis is applied to two metrics of racial discrimination: the outcome test (pierson2020large) and bounds on the ATE (knox2020administrative). There are many other metrics which have been used to determine racial bias in traffic stops and could be assessed using a similar sensitivity analysis, for example, the KPT test (knowles2001), the Threshold test (simoiu2017; pierson2018; pierson2020large), Internal Benchmarking (ridgeway2010methods), or the Hit Rates test (persico2006generalising).
Note that there are methods, such as the Veil of Darkness (grogger2006testing), which have been able to directly address the problem of data reporting rates that vary by race through modeling. In particular, their model uses ratios of probabilities (of data reporting) that can vary by race but do not vary given race and time. Not every model, however, can easily integrate probabilities of reporting, and we believe that our method of running a sensitivity analysis can benefit many models which measure racial disparities in policing.
Additionally, the methods provided here are applied exclusively to traffic stop data. However, the ideas could be expanded to data of many different types. As mentioned in Section 2, we see the importance of addressing missingness in healthcare (kaambwa2012; ward2020; hunt2021; zhang2023machine), natural disasters (jones2022human), psychology (gomer2023realistic), agriculture (robbins2013imputation), epidemiology (allotey2019multiple; balzer2020; tsiampalis2020missing), and crime (edwards2017effect; blasco2021missing; stockton2024now). A sensitivity analysis using the steps described in Figure 1 could be valuable across these and many other disciplines in which missing data convolutes the resulting analysis conclusions.
We recognize the inherent difficulty in assessing racial discrimination using observational studies with missing or poorly recorded data. We have provided a method for assessing the sensitivity of results which aim to measure racial disparities. In our own work we find that there is quite a bit of missing data in the SOPP which is not MCAR. We were able to find a substantial shift toward discrimination in considering the outcome test in Ohio. The sensitivity analysis on the ATE bounds show that there is anti-Black discrimination in multiple jurisdictions, even considering the missing race information.
Disclosure statement
The authors have no conflicts of interest to declare.
Acknowledgments
S.K. was supported in part by grants from the Pomona College SURP program, the Kenneth Cooke Summer Research Fellowship, and NSF Conference grant 2328509. Amber Lee was important to much of the work, and we appreciate her early contributions. Additionally, we appreciate the ideas put forth by Ghassan Sarkis and the following students, who worked on aspects of the project: Oliver Chang, Emma Godfrey, Will Gray, Ethan Ong, Kyle Torres, Arm Wonghirundacha, and Ivy Yuan.
Contributions
S.K. conducted the computational study and wrote the related aspects of the manuscript. J.H. conceived of the investigation, supported the computational work, and wrote sections of the manuscript.
Appendix A SOPP Data Summary
Table 4: The 20 most commonly recorded variables in the SOPP.
Covariate Number of datasets
date 88
subject race 77
outcome 74
location 73
time 72
subject sex 71
citation issued 68
subject age 53
latitude 51
longitude 51
warning issued 49
arrest made 48
search conducted 45
violation 45
officer id hash 43
contraband found 36
search basis 34
reason for stop 33
county name 31
vehicle make 29
Table 5: SOPP per variable missingness rates. List of the 88 datasets in the SOPP. Eight bold datasets were selected based on the criteria discussed in Section 3.4 NA indicates that the variable was not recorded.
Dataset Date_ range # Rows # Covariates subject_ race search_ conducted contraband_ found
AR, Little Rock 2017 - 2017 13641 13 0.051 NA NA
AZ, Gilbert 2008 - 2018 480599 12 NA NA NA
AZ, Mesa 2014 - 2019 157015 20 0.000 NA NA
AZ, Statewide 2009 - 2017 3498159 29 0.116 0.000 0.000
CA, Anaheim 2012 - 2017 87876 4 NA NA NA
CA, Bakersfield 2008 - 2018 189685 18 0.441 NA NA
CA, Long Beach 2008 - 2017 365924 27 0.000 NA NA
CA, Los Angeles 2010 - 2018 5418402 10 0.000 NA NA
CA, Oakland 2013 - 2017 133407 28 0.000 0.000 0.000
CA, San Bernardino 2011 - 2017 90523 12 NA NA NA
CA, San Diego 2014 - 2017 383027 21 0.322 0.000 0.000
CA, San Francisco 2007 - 2016 905070 22 0.000 0.000 0.000
CA, San Jose 2013 - 2018 152834 20 1.782 1.618 0.000
CA, Santa Ana 2014 - 2018 46268 15 0.225 NA NA
CA, Statewide 2009 - 2016 31778515 22 0.000 0.000 95.692
CA, Stockton 2012 - 2016 41629 17 0.495 0.000 NA
CO, Aurora 2012 - 2020 257606 16 0.146 NA NA
CO, Denver 2010 - 2018 1870731 14 NA NA NA
CO, Statewide 2010 - 2017 3112853 20 12.903 7.424 0.000
CT, Hartford 2013 - 2016 18439 26 0.000 0.000 0.058
CT, Statewide 2013 - 2015 1175339 26 0.000 0.000 0.000
FL, Saint Petersburg 2010 - 2010 16383 9 NA NA NA
FL, Statewide 2010 - 2018 7297538 34 0.025 31.516 NA
FL, Tampa 1973 - 2018 2818240 12 0.000 NA NA
GA, Statewide 2012 - 2016 1906772 19 47.112 NA NA
IA, Statewide 2006 - 2016 2441335 20 69.527 NA NA
ID, Idaho Falls 2008 - 2016 104814 13 NA NA NA
IL, Chicago 2012 - 2020 2108098 35 31.343 0.000 0.000
IL, Statewide 2012 - 2017 12748173 29 0.006 0.132 0.802
IN, Fort Wayne 2007 - 2017 265019 12 NA NA NA
KS, Wichita 2006 - 2020 1030376 22 6.535 NA NA
KY, Louisville 2015 - 2020 146562 33 0.058 0.000 NA
KY, Owensboro 2015 - 2017 6921 18 0.257 NA NA
LA, New Orleans 2010 - 2018 512092 32 3.339 0.000 0.000
MA, Statewide 2007 - 2015 3416238 24 0.049 0.000 0.000
MD, Baltimore 2011 - 2017 854759 9 NA NA NA
MD, Statewide 2007 - 2014 3669665 28 0.380 0.000 17.976
MI, Statewide 2001 - 2016 800302 20 2.185 NA NA
MN, Saint Paul 2001 - 2016 675156 16 17.477 0.000 NA
MO, Statewide 2010 - 2015 10038706 9 0.000 0.000 0.000
MS, Statewide 2013 - 2016 758412 13 0.055 NA NA
MT, Statewide 2009 - 2017 921228 30 0.012 0.000 NA
NC, Charlotte 2000 - 2015 1598453 29 0.000 0.000 0.000
NC, Durham 2001 - 2015 326024 29 0.000 0.000 0.000
NC, Fayetteville 2000 - 2015 486998 29 0.000 0.000 0.000
NC, Greensboro 2000 - 2015 600031 29 0.001 0.000 0.000
NC, Raleigh 2002 - 2015 856400 29 0.000 0.000 0.000
NC, Statewide 2000 - 2015 20286645 29 0.000 0.000 0.000
NC, Winston-Salem 2000 - 2015 452560 29 0.000 0.000 0.000
ND, Grand Forks 2007 - 2016 82535 14 1.645 NA NA
ND, Statewide 2010 - 2015 330132 12 0.037 NA NA
NE, Statewide 2002 - 2016 9031494 10 0.000 0.000 NA
NH, Statewide 2014 - 2015 259822 17 36.077 NA NA
NJ, Camden 2013 - 2020 195298 23 0.838 NA NA
NJ, Statewide 2009 - 2016 3845334 24 96.142 96.306 4.435
NV, Henderson 2011 - 2019 111090 19 2.574 NA NA
NV, Statewide 2012 - 2016 737285 11 0.068 NA NA
NY, Albany 2008 - 2017 74093 16 32.251 NA NA
NY, Statewide 2010 - 2017 7962169 19 0.000 NA NA
OH, Cincinnati 2009 - 2018 315281 25 0.000 NA NA
OH, Columbus 2012 - 2016 128157 18 0.000 0.000 NA
OH, Statewide 2010 - 2017 7225577 24 8.704 0.000 80.163
OK, Oklahoma City 2011 - 2020 945107 27 0.281 NA NA
OK, Tulsa 2009 - 2016 438248 19 0.974 NA NA
OR, Statewide 2010 - 2014 1143017 5 0.000 NA NA
PA, Philadelphia 2014 - 2018 1865096 22 0.000 0.000 0.000
RI, Statewide 2005 - 2015 509681 31 5.704 0.000 0.000
SC, Statewide 2005 - 2016 8983810 34 0.000 0.000 0.000
SD, Statewide 2012 - 2016 435895 16 NA NA NA
TN, Nashville 2010 - 2019 3092351 42 0.060 0.001 0.000
TN, Statewide 1971 - 2016 3829082 20 0.798 NA NA
TX, Arlington 2016 - 2016 112526 19 0.000 0.000 NA
TX, Austin 2006 - 2016 483255 29 0.000 0.000 0.000
TX, Garland 2012 - 2019 159845 23 0.014 NA NA
TX, Houston 2014 - 2020 2045972 21 20.168 NA NA
TX, Lubbock 2008 - 2018 546101 10 NA NA NA
TX, Plano 2012 - 2015 249043 36 0.000 0.049 0.000
TX, San Antonio 2012 - 2020 1301103 33 0.037 0.000 0.000
TX, Statewide 2006 - 2017 27426840 34 0.001 8.018 0.000
VA, Statewide 2006 - 2016 5006847 11 0.000 0.000 NA
VT, Burlington 2012 - 2020 36845 26 4.627 0.000 0.000
VT, Statewide 2010 - 2015 283285 25 1.406 0.000 0.000
WA, Seattle 2006 - 2015 319959 27 99.896 NA NA
WA, Statewide 2009 - 2018 11333425 30 28.176 0.000 0.000
WA, Tacoma 2007 - 2017 271912 15 NA NA NA
WI, Madison 2007 - 2020 331450 25 1.122 NA NA
WI, Statewide 2010 - 2016 1058902 44 14.505 0.000 0.000
WY, Statewide 2011 - 2012 173311 16 0.294 NA NA
Appendix B Causal DAG
The DAG in Figure 3 illustrates missing data as well as missing potential outcomes mechanisms in the trafficstop dataset. Here we provide detailed information about the nodes in the DAG. Additionally, Table 6 provides examples for causal relationships and missingness mechanisms that might exist in the DAG in Figure 3. We use $X$ to represent covariates. $X_{1}$ refers to completely observed covariates in the dataset, such as date, time, location, county, and police department. These variables are complete, i.e., 0% missingness, across a particular state patrol agency or municipal police department dataset in SOPP. $X_{2}$ comprises variables with missing data - subject race, subject age, stop outcome, etc. The treatment, $M$, represents whether the individual was stopped by the police. As a mediating variable, $M$ may be affected by covariates $X$ (both complete and incomplete) such as date or subject race. The outcome, $Y$, represents any post-stop outcome such as whether the individual was searched, found carrying contraband, or arrested. Unmeasured confounders, $U$ (e.g., the officer’s suspicion or their perception of the individual’s race), might affect the treatment, $M$ (whether the individual is stopped), and the outcome, $Y$ (whether they are searched).
We introduce three indicator variables $R_{X_{2}}$, $R_{M}$, and $R_{Y}$ to denote whether an instance contains missing data for partially-observed covariates ($X_{2}$), the treatment ($M$), or outcome ($Y$). moreno2018canonical also include unmeasured causes of missingness ($W$) that might affect the probability of missingness. $W$ could include factors such as the weather or departmental policy changes for the precinct that might affect data records.
Having no arrows to the missingness indicators, $M$, would mean the missingness mechanism is MCAR and entirely dictated by unmeasured factors, $W$. However, in the case of SOPP data, we notice that complete and missing covariates have the potential to affect the missingness of the stop, the outcome, and the covariates themselves as discussed in Section 3. For example, a certain police department might have higher rates of missing data in their reporting (pierson2020large) or officers might omit data at differing rates for different races (chanin2021). Hence, Figure 3 contains arrows from $X_{1}$ and $X_{2}$ to the three missingness indicators.
Additionally, policing datasets such as SOPP inherently only include observations with one level of the treatment - individuals who were stopped ($M=1$) - and contain no report on the number of individuals who were observed but not stopped ($M=0$) (knox2020administrative).
Lastly, the stop outcome, $Y$, may affect its own missingness, $M_{Y}$, as well as records of covariates, $M_{Z_{2}}$, such as subject race (i.e., officers might omit race if the individual was not searched). These causal effects are depicted in the DAG.
Table 6: Example causal relationships and missingness mechanisms to describe the DAG in Figure 3.
Arrow to
$\mathbf{X_{1}}$ $\mathbf{X_{2}}$ $\mathbf{M}$ $\mathbf{Y}$ $\mathbf{U}$ $\mathbf{R_{X_{2}}}$ $\mathbf{R_{M}}$ $\mathbf{R_{Y}}$ $\mathbf{W}$
Arrow from $\mathbf{X_{1}}$ More stops in some locations More searches in some locations More race missingness early morning Less stops early morning Less searches early morning
$\mathbf{X_{2}}$ More stops among Black civilians More searches among Black civilians More race missingness for Black stops (chanin2021) Less stops among white civilians Complacency in search records for some race
$\mathbf{M}$ If someone is not stopped, they cannot be searched If someone is not stopped, their race records are missing If someone is not stopped, their stop records are missing
$\mathbf{Y}$ Complacency in race records if not searched Complacency in stop records if not searched Complacency in search records if not searched
$\mathbf{U}$ Police suspicion Police suspicion
$\mathbf{R_{X_{2}}}$
$\mathbf{R_{M}}$
$\mathbf{R_{Y}}$
$\mathbf{W}$ More race missingness in some weather Less stops in some weather More search missingness in some weather
Appendix C Outcome test: additional results
Figure 11: Sensitivity analysis for 32 counties in Colorado. Open black circles are sized by number of searches.
Figure 12: Sensitivity analysis for 51 counties in Wisconsin. Open black circles are sized by number of searches.
Figure 13: Sensitivity analysis for 31 counties in Washington. Open black circles are sized by number of searches.
Figure 14: Sensitivity analysis for 60 police and sheriff departments in Maryland. Open black circles are sized by number of searches.
Figure 15: Sensitivity analysis for 87 counties in Ohio. Open black circles are sized by number of searches.
Appendix D ATE Bounds: additional results
D.1 Random Assignment
Figure 16: Dashed line at zero indicates no differential racial impact. Disparity is the difference in proportion of search for Black minus proportion of search for white. Disparity is measured as bounds on the search ATE, given recorded outcome, for different levels of the $\rho$ parameter (the overall proportion of racially discriminatory stops). The $x$-axis of each plot presents a different allocation of assigning NA race observations to either Black or white.
Figure 17: Dashed line at zero indicates no differential racial impact. Disparity is the difference in proportion of search for Black minus proportion of search for white. Disparity is measured as bounds on the search ATE, given recorded outcome, for different levels of the $\rho$ parameter (the overall proportion of racially discriminatory stops). The $x$-axis of each plot presents a different allocation of assigning NA race observations to either Black or white.
Figure 18: Dashed line at zero indicates no differential racial impact. Disparity is the difference in proportion of search for Black minus proportion of search for white. Disparity is measured as bounds on the search ATE, given recorded outcome, for different levels of the $\rho$ parameter (the overall proportion of racially discriminatory stops). The $x$-axis of each plot presents a different allocation of assigning NA race observations to either Black or white.
D.2 Extreme Cases
Figure 19: Dashed line at zero indicates no differential racial impact. Disparity is the difference in proportion of search for Black minus proportion of search for white. Disparity is measured as bounds on the search ATE, given recorded outcome, for different levels of the $\rho$ parameter (the overall proportion of racially discriminatory stops). The $x$-axis of each plot presents the ignore NA case as well as the two most extreme scenarios of allocating NA race observations to either Black or white. Note that for CO, Statewide, the disparity estimates can change direction from discrimination against Black drivers to the opposite direction.
Figure 20: Dashed line at zero indicates no differential racial impact. Disparity is the difference in proportion of search for Black minus proportion of search for white. Disparity is measured as bounds on the search ATE, given recorded outcome, for different levels of the $\rho$ parameter (the overall proportion of racially discriminatory stops). The $x$-axis of each plot presents the ignore NA case as well as the two most extreme scenarios of allocating NA race observations to either Black or white. Note that for MD, Statewide, the ATE bounds are always above zero, even in the most extreme cases, indicating racial discrimination against Black drivers.
Figure 21: Dashed line at zero indicates no differential racial impact. Disparity is the difference in proportion of search for Black minus proportion of search for white. Disparity is measured as bounds on the search ATE, given recorded outcome, for different levels of the $\rho$ parameter (the overall proportion of racially discriminatory stops). The $x$-axis of each plot presents the ignore NA case as well as the two most extreme scenarios of allocating NA race observations to either Black or white. Note that for WI, Statewide, the ATE bounds are almost always above zero, indicating discrimination against Black drivers.
Appendix E Data Repository/Code
The Open Policing Project datasets can be found at https://openpolicing.stanford.edu/data/. Code to replicate our findings is available at https://github.com/saatvikkher/missing-traffic.